OpenStudy (anonymous):

solve \[\log_{7}(x+1)+\log_{7}(x-5)=1 \]

OpenStudy (ghazi):

use base change formula

OpenStudy (ghazi):

\[\log _{7}(x+5)=\frac{ \log (x+5) }{ \log 7 }\]

OpenStudy (ghazi):

after that use\[\log (a*b)= \log a + \log b\]

OpenStudy (ghazi):

@alexeis_nicole can you do it?

OpenStudy (anonymous):

its \[\log_{7}(x-5) .. \] not plus :S

OpenStudy (ghazi):

sorry, somewhat it will be like \[\log (\frac{ x+1 }{ 7 }*\frac{ x-5 }{ 7 })\]

OpenStudy (ghazi):

can you do it now? take the log on the other side it will be e^1=e=2.732 and you will get a quadratic equation, hope you can do it now

OpenStudy (anonymous):

i got to log[(x^2-4x-5)/7]=1

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

now, \[\frac{ (x+1)(x-5) }{ 49 }=e^1=e=2.73\] then multiply both sides by 49

OpenStudy (anonymous):

uhh im kinda lost now.. /:the way my book shows it is different from how you are @ghazi

OpenStudy (ghazi):

well you just need to use properties of logs that is it or may be you can replace e by 10 if the base of log is 10

OpenStudy (ghazi):

\[\frac{ (x+1)(x-5) }{ 49 }=10\] then \[(x+1)(x-5)=490\]

OpenStudy (anonymous):

\[\log_{7}((x+1)(x-5))=1\] (x+1)(x-5)=7^1 x^2-4x-5=7 x^2-4x-12=0 (x-6)(x+2) x=6 x=-2